OpenStudy (anonymous):

(g ◦ f)^-1 g and f are functions. could this be rewritten as g^-1 ◦ f^-1?

jimthompson5910 (jim_thompson5910):

h(x) = f o g h(x) = f(g(x)) y = f(g(x)) x = f(g(y)) f^(-1)(h(x)) = g(y) g^(-1)(f^(-1)(h(x))) = y y = g^(-1)(f^(-1)(h(x))) h^(-1)(x) = g^(-1)(f^(-1)(h(x))) h^(-1)(x) = g^(-1) o f^(-1)

jimthompson5910 (jim_thompson5910):

So if h(x) = f o g then h^(-1)(x) = g^(-1) o f^(-1)

jimthompson5910 (jim_thompson5910):

notice the flip

OpenStudy (anonymous):

functions, numbers, elements of a group, it makes no difference \[(ab)^{-1}=b^{-1}a^{-1}\] and the check is that \[abb^{-1}a^{-1}=aa^{-1}=I\]

OpenStudy (anonymous):

Thanks guys I've got this question Consider the functions given by f = {(1, 1),(2, 1),(3, 1),(4, 3),(5, 2)} from {1, 2, 3, 4, 5} to {1, 2, 3, 4} g = {(1, 2),(2, 1),(3, 4),(4, 3)} from {1, 2, 3, 4} to {1, 2, 3, 4, 5} and I need to find (g ◦ f)^-1 and say if its a function or not so is it correct to write this as g^-1 (f^-1)?

jimthompson5910 (jim_thompson5910):

(g ◦ f)^-1 = f^-1 o g^-1 using the idea I showed above

jimthompson5910 (jim_thompson5910):

notice this: say you plug in x = 2 into g^-1 the output will be g^(-1)(2) = 1 using the table above So f^-1 o g^-1 will turn into f^-1(1) but f^-1(1) = {1, 2, 3} which means that (g ◦ f)^-1 = f^-1 o g^-1 is NOT a function

OpenStudy (anonymous):

Thank you so much for your help

OpenStudy (anonymous):

are you sure this is correct? why is f^-1(1) = {1, 2, 3} not correct? surely this is like having multiple roots in the equation

jimthompson5910 (jim_thompson5910):

notice how when x = 1, f(x) = 1 when x = 2, f(x) = 1 So this suggests that f is not one-to-one

jimthompson5910 (jim_thompson5910):

which means that the inverse of f is not a function

OpenStudy (anonymous):

so if i plug g^-1 into f^-1 i get = {(1,1), (1,2), (3,3), (1,4)} this is not a function but how do i phrase it properly because we have to explain why.

jimthompson5910 (jim_thompson5910):

if x = 2, then plugging x = 2 into the inverse of g, you get 1 do you see this?

OpenStudy (anonymous):

yes I see what you are saying so is writing (g ◦ f)^-1(1) = {1, 2, 4} good enough of an explanation?

jimthompson5910 (jim_thompson5910):

yes, you can say that because one input gives you multiple outputs means it's not a function

OpenStudy (anonymous):

when you said 'So f^-1 o g^-1 will turn into f^-1(1)' but in here you could also have 2 or 3 not just 1.

jimthompson5910 (jim_thompson5910):

no, you could have 2 or 3 at the end f^-1(1) is just an intermediate step

OpenStudy (anonymous):

the inverse is f^-1 = {(1, 1),(1, 2),(1, 3),(4, 3),(5, 2)} so if you plug 1 into here you get multiple results

jimthompson5910 (jim_thompson5910):

exactly

jimthompson5910 (jim_thompson5910):

but my step f^-1(1) is the step right before you get those multiple outputs